OpenStudy (lgbasallote):

i got stuck lol... \(\Huge \int \frac{\cos x dx}{(\sin x)^3 + \sin x}\)

OpenStudy (anonymous):

Take sin x =t dt = cosx dx \[\int\limits_{}^{}dt/ t(1+t^2)\] Apply partial fraction. . I hope you know what is partal fraction

OpenStudy (anonymous):

\[\int \frac{du}{u^3+u}\]

OpenStudy (lgbasallote):

ahh partial fractions of course...but for the sake of it..would you mind showing it?

OpenStudy (lgbasallote):

just a wild question...is trig sub applicable in u^2 + u?

OpenStudy (anonymous):

\[\frac{a}{x}+\frac{bx+c}{x^2+1}=\frac{1}{x(x^2+1)}\] \[a(x^2+1)+x(bx+c)=1\] let \(x=0\) get \(a=1\)

OpenStudy (anonymous):

In any case. We can write \[1/(t)(1+t^2) = A/t + (Bx+C)/1+t^2\] \[1=A(1+t^2) +(Bx+C)(t)\] Where A,B,C are constants/ Now substitute different values of x to get A, B and C Finally integrate :)

OpenStudy (anonymous):

and since \[a(x^2+1)+x(bx+c)=1\] \[x^2+1+x(bx+c)=1\] \[x^2+bx^2=(1+b)x^2\implies b=-1\] \]

OpenStudy (lgbasallote):

how did c become zero?

OpenStudy (anonymous):

and \(cx\implies c=0\) s you get \[\int\frac{du}{u}-\int \frac{udu}{u^2+1}\]

OpenStudy (anonymous):

c is zero because if you expand the right left hand side you end up with a \(cx\) term, but on the right hand side there is no x term. that means \(c=0\)

OpenStudy (lgbasallote):

ahh i see..silly me =))

OpenStudy (lgbasallote):

so from that im seeing ln u - [ln (u^2 + 1)]/2

OpenStudy (anonymous):

just a wild question...is trig sub applicable in u^2 + u? Yes. I think this method will be faster/ \[\int\limits_{}^{}dt/(t)(1+t^2)\] Substitute t=tanx \[\int\limits_{}^{} \sec^2x dx/ \tan x(1+\tan^2x) \] Since 1+tan^2 (x) = sec^2 (x) Therefore sec^2(x) in numerator and denominator gets cancelled. So we are left with \[\int\limits_{}^{}\cot x dx = \log \left| sinx \right| + c\]

OpenStudy (anonymous):

Short and sweet :)

OpenStudy (anonymous):

Final answer is \[\log \left| \sin(\tan^{-1} t) \right|+C\]

OpenStudy (lgbasallote):

but wolfram gave a different answer :/

OpenStudy (lgbasallote):

i got ln (sinx) - 1/2 ln (sin^2 x + 1)

OpenStudy (anonymous):

We always have many possible answers to the same integration question. They all differ by constant. can you point out any mistake in my method ??

OpenStudy (lgbasallote):

well in the trig substitution we were the same until log |sin x| i think what x is...we're different haha

OpenStudy (lgbasallote):

no..we're the same all through out...in log|sin theta| you shouldve gotten log|(t)/sqrt(t^2 + 1))

OpenStudy (lgbasallote):

log t - 1/2 log | t^2 + 1| change back to sinx log |sin x| - 1/2 log{sin^2 x + 1|

OpenStudy (lgbasallote):

you messed up when you reused x :P when it was in the original problem already

OpenStudy (anonymous):

Yes. I agree @lgbasallote . YOu are right I messed it up taking x

OpenStudy (lgbasallote):

though what i dont get is why the second term of wolfram is -1/2 ln (3 - cos(2x)) can you provide me with an answer??

OpenStudy (anonymous):

ok. Which method-- partial fraction or the trig substitution?

OpenStudy (lgbasallote):

partial

OpenStudy (anonymous):

ok. :) GIve me 5 minutes. I am very slow in typing

OpenStudy (anonymous):

\[I= \int\limits_{}^{} dt/t - \int\limits_{}^{} tdt/1+t^2\] Let p=1+t^2 dp/2 = t* dt \[I=\log t - (1/2) \int\log dp/p\] \[I=\log t -(1/2) \log p + C\] substituting p= 1+t^2 \[I=\log t- (1/2)(\log (1+t^2)) +C\] substituting t= sin x \[I = \log (\sin x) - (1/2)(\log(1+(\sin x)^2)) +C\]

OpenStudy (anonymous):

Now We know that cos 2x = 1 - 2 (sinx)^2 Therefore (sinx)^2 = (1 - cos 2x)/2

OpenStudy (anonymous):

\[I = \log (\sin x) - (1/2)(\log(1+(1-\cos 2x)/2)\] \[I = \log (\sin x) - (1/2)(\log(2+1-\cos 2x)/2)\] \[I = \log (\sin x) - (1/2)(\log((3-\cos 2x)/2))\]

OpenStudy (anonymous):

@lgbasallote , Hope you got it :)

OpenStudy (anonymous):

lol I took 20 minutes to type this. My bad :(

OpenStudy (anonymous):

@Callisto , beat this record. 20 minutes to type half solution to a integral problem :)

OpenStudy (callisto):

@shivam_bhalla lol I'll try if I have 30 minutes :P Perhaps in your question ? lol

OpenStudy (lgbasallote):

lol it seems they were just the same :P wolfram just did some crazy thing..20 mins wasted >:)))

OpenStudy (anonymous):

Seriously :P I told you. Most of the answers are simply manipulating among trig identities or differing by a constant

OpenStudy (callisto):

But I always don't know which trigo identities to use :(

OpenStudy (anonymous):

@Callisto , it only comes by practice and the experience gained out of practicing :)

OpenStudy (callisto):

That's why I'm trying to remember all the cases :(